The charge C, for renting an air hammer from Taylor and Son Equipment Rental is a linear function the number of days (n) in the rental period. The charge is
$113 for 2 days and $209 for 5 days.
(a) Write C as a linear function of n. Your equation should look like C= an + b
Where you need to find the values of a + b using the information above as
ordered pairs (2, 113) and (5, 209)

(b) What would the charge be for 4 days

Answer by TakeishaM(13)

First, find the slope using the points. Slope (M) = y2-y1/x2-x1
(2,113) (5,209)
M = (209-113)/(5-2)
= 96/3
= 32
Now pick a point, say (2,113). Plug in the x,y and m values in the equation and solve for b
113 = 32(2) + b
113 = 64 + b
113-64 = 64-64+b
49 = b
Your equation is C = 32n + 49
The cost for 4 days...just plug in 4 for n. Multiply 4 times 32 and add that to 49
C = 32(4)+49
C = 177